OpenStudy (anonymous):

Six lines are drawn through a circle. What is the maximum number of regions into which the lines can divide the interior of the circle?

OpenStudy (anonymous):

Not sure, but I have 12

OpenStudy (anonymous):

Sorry I'm late, but apparently the answer is 22.

OpenStudy (anonymous):

Nevermind. I figured it out. It's like the fibonacci sequence. 1 + 1 = 2 2 + 2 = 4 3 + 4 = 7 4 + 7 = 11 5 + 11 = 16 6 + 16 = 22 That's how they figured it out...